3x-8y=11
Solve for y.

Answer by harpazo(655)

3x-8y=11
Solve for y.
3x-8y = 11
We need to get y alone on one side of the equation.
Subtract 3x from both sides.

-8y = -3x + 11
Divide both sides by -8.
y = (-3x+ 11)/-8

y = (3/8)(x) - 11/8